# Black holes in the expanding Universe

### v2

@@ Introduction.
The relation (17), regularizing the pressure and curvature for the metric (1) at the event horizon of a black hole, agrees with the actual formation of an event horizon after an infinite cosmic time $\tau$ because of gravitational time dilation Tolman ; LL2 . As $\tau\to\infty$, the Friedmann equations (3) in the presence of a positive cosmological constant give $H\to(\Lambda/3)^{1/2}$ (17). Without the cosmological constant, $H\to 0$. In both cases, $\dot{H}\to 0$ as $\tau\to\infty$, so the regularization of an event horizon occurs in the same time limit as its formation Kaloper . However, the metric (1) describes a black hole that has already formed. Therefore, to regularize the horizon, one must use the condition (17) instead of the limit $H\to H_{\textrm{hor}}$ as $\tau\to\infty$. A more complete consideration should use the comoving coordinates Tolman to describe gravitational collapse in an expanding universe condensation .
More generally, the condition $F/r+F^{\prime}=0$ is satisfied for $F=A(\tau)/r$, where $A(\tau)$ is some function of the time coordinate. Consequently, $\dot{a}/a=H(r)$ near the horizon, which gives $d(\ln a)/d\tau=H(r)$ and thus $\ln a=H(r)\tau+\ln B(r)$, where $B(r)$ is some function of the radial coordinate. Substituting the resulting $a(\tau,r)=B(r)e^{H(r)\tau}$ into $F=a^{\prime}/a$ gives $A(\tau)-rB^{\prime}(r)/B(r)=r\tau H^{\prime}(r)$, which gives $H=\textrm{const}$, $A=\textrm{const}$, and $B(r)=Cr^{A}$. The first equation in (14) gives $\kappa\epsilon_{\textrm{hor}}=3H^{2}_{\textrm{hor}}-[(1/16a^{2}r^{2})(2A+A^{2})]_{\textrm{hor}}$. The above relations give the dependence of the scale factor on $\tau$ and $r$. This solution represents a black hole that is accreting matter. In this case, the constant $H_{\textrm{hor}}$ is different from (17) and depends on the dynamics of accretion.
The spacetime of a black hole with or without accretion, embedded in the flat, expanding Universe, is determined by the field equations (14) with the boundary conditions (15) and (16). These conditions, regularizing the spacetime at the horizon, are possible because the scale factor is inhomogeneous and depends on $r$.
Effect of universe expansion on an event horizon.
If the Hubble parameter near an event horizon is constant and equal to (17), then $F_{\textrm{hor}}=0$, so the scale factor near the horizon is almost homogeneous: $a(\tau)$. In this case, the following coordinate transformations Robertson :
$\rho=a(\tau)r(1+M)^{2},\quad t=\tau+8(Gm)^{2}\sqrt{\frac{\Lambda}{3}}F\Bigl{(}\frac{1-2Mr}{1+2Mr}\Bigr{)},$ (18)
where the function $F$ is given by
$F(x)=\int\frac{dx}{(1-x^{2})[x^{2}(1-x^{2})^{2}-(4/3)(Gm)^{2}\Lambda]},$ (19)
transform the metric (1) into the Kottler (Schwarzschild–de Sitter) metric Kottler :
$ds^{2}=\Bigl{(}1-\frac{2Gm}{\rho}-\frac{1}{3}\Lambda\rho^{2}\Bigr{)}dt^{2}-\frac{d\rho^{2}}{1-2Gm/\rho-\Lambda\rho^{2}/3}-\rho^{2}d\theta^{2}-\rho^{2}\sin^{2}\theta\,d\phi^{2}.$ (20)
@@
The Kottler metric describes a centrally symmetric black hole in a spacetime with a cosmological constant, reducing to the de Sitter metric deSitter in the absence of the black hole and to the Schwarzschild metric in the absence of the cosmological constant. Consequently, the radius $\rho_{\textrm{hor}}$ of a black hole, given by the condition $01-2Gm/\rho_{\textrm{hor}}-\Lambda\rho^{2}_{\textrm{hor}}/3=$ 0, is constant in the absence of accretion. A black hole does not grow with the expansion of the universe. A constant Hubble parameter at an event horizon means that points in space near the horizon move away from it (space expands), but the horizon remains constant in time. For a black hole accreting matter, $F_{\textrm{hor}}\neq 0$, and transformations analogous to (18) would give a metric with a mass increasing in time.
Conclusion.
@@
The McVittie metric does not describe a physical black hole in an expanding universe because the curvature scalar and pressure at the event horizon are infinite. We extended this metric to an inhomogeneous scale factor, which depends on both the time and radial coordinate. We showed that the resulting energy density is inhomogeneous, and the pressure is inhomogeneous and anisotropic, but far from the black hole they tend to homogeneous and isotropic functions of time, in accordance with the observed Universe on the large scale. We showed that the inhomogeneity of the scale factor eliminates those infinities by imposing the conditions (15) and (16) at the horizon. Consequently, the Hubble parameter at the event horizon of every centrally symmetric black hole in the absence of accreting matter is equal to the same constant, determined by the cosmological constant (including zero) in the relation (17).
@@
We also showed that, because of the constancy of the Hubble parameter at an event horizon and the resulting equivalence of the generalized McVittie metric and the Kottler metric, black holes do not grow with the universe expansion. This result is contrary to a hypothesis in growth that black holes may be growing as the universe expands, which aimed to explain how the enormous supermassive black holes at the centers of most galaxies have formed, and agrees with the results of Rudeep . Black holes can grow only if they accrete matter or merge with other black holes. We showed that the generalized McVittie metric with an inhomogeneous scale factor could also describe black holes accreting matter, embedded within galaxies and other large-scale structure instead of a region of the FLRW spacetime.
@@
Finally, the results of this work are also valid in the Einstein–Cartan theory of gravity EC , which removes the symmetry constraint on the affine connection and relates its antisymmetric part, the torsion tensor Niko , to the intrinsic angular momentum (spin) of the matter composed of fermions, described by the Dirac equation. In the presence of torsion, which manifests at extremely high densities as a repulsive force, gravitational singularities in black holes may be eliminated, and the singular big bang may be replaced by a regular big bounce universe . In vacuum, torsion vanishes and this theory reduces to general relativity, passing all its observational tests and giving the regularizing conditions (15) and (16) at an event horizon.
@@
I am grateful to Francisco Guedes and my Parents, Bożenna Popławska and Janusz Popławski, for supporting this work.

### v3

@@ 8. Constancy of Hubble parameter at an event horizon
This result equates two boundary values of $H$ at different ends of the interval: an event horizon $\rho=\rho_{\textrm{hor}}$ and a future infinity $\tau\to\infty$. It is consistent with the actual formation of an event horizon after an infinite cosmic time $\tau$ because of gravitational time dilation Tolman ; LL2 . As $\tau\to\infty$, the Friedmann equations (3) in the presence of a positive cosmological constant give $H\to(\Lambda/3)^{1/2}$. Without the cosmological constant, $H\to 0$. In both cases, $\dot{H}\to 0$ as $\tau\to\infty$, so the regularization of an event horizon occurs in the same time limit as its formation Kaloper . A more complete consideration should use the comoving coordinates Tolman to describe gravitational collapse, forming a black hole, in an expanding Universe condensation .
9. Effect of Universe expansion on an event horizon
@@
At the event horizon of a black hole, described by the metric (16), the Hubble parameter is constant and equal to (34). Consequently, in the infinitesimal vicinity of such a black hole, where the scale factor does not vary significantly with $\rho$, the transformations (4) and (6) turn the metric (16) into the Kottler metric (7). Therefore, a black hole does not grow with the expansion of the Universe; it can grow only as a result of accretion. Points in the space near the horizon move away from it, as the space expands, but the size of the horizon $r_{\textrm{hor}}$ remains constant in time.
10. Conclusion
@@
The McVittie metric does not describe a physical black hole in an expanding Universe because the curvature scalar and pressure at the event horizon are infinite. We extended this metric to a metric with an inhomogeneous scale factor, which depends on both the time and radial coordinate (16). The resulting energy density is inhomogeneous, and the pressure is inhomogeneous and anisotropic. We showed that the inhomogeneity of the scale factor eliminates those infinities by imposing the conditions (25) and (26) at the horizon. We considered a special case of the metric, for which the second condition is satisfied also outside the horizon. We found the corresponding form of the scale factor (27), satisfying the regularity condition (30), and a condition under which the energy density and pressure far from the black hole tend to nearly homogeneous and isotropic functions of time, in accordance with the observed Universe on the large scale.
@@
We showed that the Hubble parameter at the event horizon of every centrally symmetric black hole is equal to the same constant, determined by the cosmological constant (including zero) in the relation (34). We also showed that, because of the resulting equivalence of the generalized McVittie metric and the Kottler metric, black holes do not grow with the Universe expansion. This result is contrary to a hypothesis in growth that black holes may be growing as the Universe expands, which aimed to explain how the enormous supermassive black holes at the centers of most galaxies have formed, and agrees with the results of Rudeep . Black holes can grow only if they accrete matter or merge with other black holes.
@@
The results of this work are also valid in the Einstein–Cartan theory of gravity EC , which removes the symmetry constraint on the affine connection and relates its antisymmetric part, the torsion tensor Niko , to the intrinsic angular momentum (spin) of the matter composed of fermions, described by the Dirac equation. In the presence of torsion, which manifests at extremely high densities as a repulsive force, gravitational singularities in black holes may be eliminated, and the singular big bang may be replaced by a regular big bounce universe . In vacuum, torsion vanishes and this theory reduces to general relativity, passing all its observational tests and giving the regularizing conditions (25) and (26) at an event horizon.
@@
I am grateful to Francisco Guedes and my Parents, Bożenna Popławska and Janusz Popławski, for supporting this work. I thank the reviewers for their helpful comments and suggestions, which have strengthened the presented analysis.